OpenStudy (jskinny0831):

What is the sum of the geometric series in which a1 = 7 , r = 3 and an = 1701? Sn = 2,548 Sn = 851 Sn = 58,824 Sn = −851

OpenStudy (anonymous):

find n by the formula \[a _{n}=a _{1}r ^{n-1}\] find n then \[S _{n}=a _{1}\frac{ r^n-1 }{ r-1 }\]

OpenStudy (anonymous):

\[1701=7*3^{n-1}\] \[3^{n-1}=\frac{ 1701 }{ 7 }=243=3^5,n-1=5,n=5+1=6\] \[S _{n}=7\frac{ 3^n-1 }{ 3-1 }=7\frac{ 3^6-1 }{ 3-1 }=7\frac{ 729-1 }{ 2 }=?\]